The pilot of the airplane which has been diving at a speed of 540 km/h pulls out of the dive following a circular path at a constant speed.
- What is the minimum radius of the planes circular path in order that the acceleration of the pilot at the lowest point not exceed 7G?
- What is the apparent weight of an 80.0kg pilot at the lowest point of the pullout.
1
Expert's answer
2019-04-23T11:03:19-0400

The centripetal acceleration

"a=\frac{v^2}{R}"

So

"R=\frac{v^2}{a}=\frac{v^2}{7g}"

"=\frac{(540\times \frac{1000\:\rm{m}}{3600\:\rm{s}})^2}{7\times 9.8\:\rm{m/s^2}}=328\:\rm{m}"

The apparent weight of an 80.0 kg pilot at the lowest point of the pullout

"P=m(g+a)=8mg=8\times 80\:\rm{kg}\times 9.8\:\rm{m/s^2}=6272\:\rm{N}"